OpenStudy (anonymous):

An aircraft carrier made a trip to Guam and back. The trip there took three hours and the trip back took four hours. It averaged 6 km/h on the return trip. Find the average speed of the trip there

OpenStudy (mathmate):

It's not a very fast plane. Can you check for typos?

OpenStudy (anonymous):

haha

OpenStudy (mathmate):

Oh, sorry, it's a carrier!

OpenStudy (mathmate):

Figure out the distance from the return trip, and divide double the distance by 7 hours.

OpenStudy (mathmate):

The one-way distance is 6km/h*4 hours = 24 km.

OpenStudy (anonymous):

Speed = Distance/Time Speed * Time = Distance => Distance is constant. Therefore S1 * T1 = S2 * T2 S1=6 T1=4 S2=? T2=3 6*4 = ?*3 24/3 = ? ? = 8km/h

OpenStudy (mathmate):

Sorry again, the average speed to go there is the distance divided by 3 hours = 24km/3 = 8 km/h.